### v1

@@ 6 Multiplier bias and uniqueness of an equilibrium
where ${\bf E}(\cdot)$ denotes the set of equilibria. This means that the set of Nash equilibria is unaffected by positive rescaling of the objective functions. An interesting aspect, however, is that except in the orthogonal case, i.e., when $\mathcal{C}$ has the form $\mathcal{C}=\mathcal{C}_{1}\times\cdots\times\mathcal{C}_{N}$, we have ${\bf VE}(G_{r})\neq{\bf VE}(G)$, where ${\bf VE}(\cdot)$ is the set of variational equilibria. We show below, extending the results of [rosen_existence_1965] to infinite dimensions, that an $r$-dependent structural condition on $G_{r}$ ensures that ${\bf VE}(G_{r})$ is a singleton. Hence, any numerical method designed to compute a variational equilibrium will converge to the same one. Since ${\bf VE}(G_{r})\subset{\bf E}(G_{r})={\bf E}(G)$ and $r\in(\mathbb{R}_{+}^{*})^{N}$ must be chosen so that the structural condition holds, we are still computing a Nash equilibrium of the original game.
@@
We begin with the non-generalized situation, i.e., we assume that ${\bf X}(x)\equiv X^{\rm{ad}}$ even if ${\bf X}_{i}$ has the form (2.2). This corresponds to the case in which $\mathcal{C}$ is a Cartesian product. The result below extends [rosen_existence_1965, Theorem 2] to the infinite-dimensional setting. In its proof we need the normal cone to $\mathcal{C}$ at some $x\in\mathcal{C}$ which is given by
@@
It follows from Lemma 4.1 and from convex optimization that if $\overline{x}\in\mathcal{C}$ is a Nash equilibrium, then there exists $\mu_{\overline{x}}\in N_{\mathcal{C}}(\overline{x})$ such that
@@
$\alpha=d(\overline{x},r)(\overline{y}-\overline{x})+d(\overline{y},r)(\overline{x}-\overline{y}),\qquad\beta=\mu_{\overline{x}}(r\cdot(\overline{y}-\overline{x}))+\mu_{\overline{y}}(r\cdot(\overline{x}-\overline{y})).$
@@
It follows from 6.3 and 6.4 that when the constraint map is constant, the variational equilibrium is unique and independent of $r$. However, in the non-constant case, although the scaling vector $r$ cannot alter the set of Nash equilibria, it may affect the set of variational equilibria – a feature of practical interest given their computational advantages over the full equilibrium set. The next proposition quantifies this effect.
@@
We do not include the proofs of 6.4 and 6.5 because the arguments follow the same structure as 6.3. The corresponding finite-dimensional proofs are found in [rosen_existence_1965].
@@
Diagonal strict convexity also implies a strong stability property, naturally formulated in terms of an associated ODE-type dynamical system that generates a trajectory of strategies. This dynamical perspective has served as a foundation for algorithms computing variational equilibria and offers a complementary viewpoint to modern numerical methods for variational inequalities; see, for example, [facchinei_generalized_2007, facchinei_decomposition_2011] and references therein. Beyond its algorithmic relevance, this interpretation provides a useful conceptual and historical connection between equilibrium theory and dynamical systems.
7 Geometric forms of generalized games
Geometric formulations of games, which are based on preference relations rather than objective functionals, are less common in the generalized Nash framework, especially in infinite dimensions. Most existing theory has been developed in analytic form, where each player solves an optimization problem and geometric structures appear only implicitly through fixed-point arguments. In this section, we present a geometric counterpart to our setting. Geometric games are interesting in their own right: they separate the modeling of preferences from numerical optimization, can be more natural in applications where utilities are only partially specified, and provide an alternative route to existence results and structural subclasses such as potential games. Our aim here is not to build a complete geometric theory matching all analytic results, but to show that many of the structural insights developed above extend naturally to preference maps and to indicate which additional tools would be needed for a fully parallel treatment.
Geometrically, the goal of each player is modeled by a preference map $\mathcal{P}_{i}:X\to 2^{X_{i}}$ defined in this way: for profile $x\in X$, $\mathcal{P}_{i}(x)$ contains those strategies $\hat{x}_{i}\in X_{i}$ that player $i$ strictly prefers over $x_{i}$ when the opponents’ strategies $x_{-i}$ are fixed. Setting $\mathcal{P}=\mathcal{P}_{1}\times\cdots\times\mathcal{P}_{N}$, we write $G=({\bf X},\mathcal{P})$ for a generalized game given in geometric form and refer to it as a geometric game.
The notion of equilibrium has a natural formulation in this setting.
Definition 7.1 (Geometric GNE).
We say that $\overline{x}\in{\bf X}(\overline{x})$ is a geometric generalized Nash equilibrium of the game $G=({\bf X},\mathcal{P})$ if
$\mathcal{P}_{i}(\overline{x})\cap{\bf X}_{i}(\overline{x}_{-i})=\emptyset,$ (7.1)
for all $i\in I$.
In words, at a geometric equilibrium no player has a strictly preferred feasible alternative strategy. The regularity and convexity properties of the preference maps, needed to obtain results similar to the ones for analytic games, are listed below.
Assumption 7.2 (Preference maps).
For each $i\in I$,
  • (i)
    the graph of $\mathcal{P}_{i}:X\to 2^{X_{i}}$ is open in $X\times X_{i}$;
  • (ii)
    for each fixed $\hat{x}_{-i}\in X_{-i}$, the set $\mathcal{P}_{i}(x_{i},\hat{x}_{-i})$ is nonempty and convex for all $x_{i}\in X_{i}$;
  • (iii)
    for each fixed $\hat{x}_{-i}\in X_{-i}$ and all $x_{i}\in X_{i}$, we have $x_{i}\in\overline{\mathcal{P}_{i}(x_{i},\hat{x}_{-i})}\setminus\mathcal{P}_{i}(x_{i},\hat{x}_{-i}).$
Recall that, for the analytic case, we investigated the role of lower semicontinuity in applying Kakutani’s fixed-point theorem. In the geometric case, the situation is different: even the construction of a best-response map relies on lower semicontinuity of the constraint maps. We briefly recall the classical existence proof for the geometric case.
Theorem 7.3.
For each $i\in I$, assume that the constraint map ${\bf X}_{i}:X_{-i}^{\rm ad}\to 2^{X_{i}^{\rm ad}}$ is lower semicontinuous. If $X$ is a separable Banach space and Assumption 7.2 holds, then the geometric GNEP $G=({\bf X},\mathcal{P})$ has an equilibrium.
Proof.
From the definition above, $x$ is a geometric generalized Nash equilibrium precisely when $x\in{\bf X}(x)$ and
$\mathcal{P}_{i}(x)\cap{\bf X}_{i}(x_{-i})=\emptyset\quad\text{for all }i\in I.$
For each $i\in I$, define
$\mathcal{C}_{i}(x)=\mathcal{P}_{i}(x)\cap{\bf X}_{i}(x_{-i}),\quad\text{and}\quad F_{i}=\{x\in X^{\rm ad}:\mathcal{C}_{i}(x)\neq\emptyset\}.$
The set $F_{i}$ consists of strategy bundles that cannot be equilibria, since for such an $x$, player $i$ has at least one strictly preferred feasible alternative. Under Assumption 7.2 and the standing assumptions on ${\bf X}_{i}$, the restricted maps $\mathcal{C}_{i}|_{F_{i}}$ are lower semicontinuous; see [michael_continuous_1956, tian_equilibrium_1990]. Moreover, because each $X_{i}^{\rm ad}$ is compact, these maps admit a continuous selection [michael_continuous_1956]. That is, there exist continuous functions
$f_{i}:F_{i}\to X_{i}^{\rm ad}\quad\text{such that}\quad f_{i}(x)\in\mathcal{C}_{i}(x)\quad\text{for all }x\in F_{i}.$
We now define, for each $i\in I$,
$\mathfrak{B}_{i}(x)=\begin{cases}f_{i}(x),&\text{if }x\in F_{i},\\[2.15277pt] {\bf X}_{i}(x_{-i}),&\text{if }x\notin F_{i},\end{cases}$
and introduce the (set-valued) best-response map
$\mathfrak{B}:X^{\rm ad}\to 2^{X^{\rm ad}},\quad\text{with }\mathfrak{B}(x)=\prod_{i=1}^{N}\mathfrak{B}_{i}(x).$
By construction, $\mathfrak{B}(x)$ is closed, convex, and compact-valued. Moreover, fixed points of $\mathfrak{B}$ coincide with geometric Nash equilibria. Indeed, suppose $\overline{x}$ is a fixed point. If $\overline{x}\in F_{i}$ for some $i$, then
$\overline{x}_{i}=f_{i}(\overline{x})\in\mathcal{C}_{i}(\overline{x})\subset\mathcal{P}_{i}(\overline{x}),$
which contradicts the equilibrium condition (7.1). Thus a fixed point cannot belong to any $F_{i}$, and therefore satisfies $\mathcal{P}_{i}(\overline{x})\cap{\bf X}_{i}(\overline{x}_{-i})=\emptyset$ for all $i$.
To apply Kakutani’s fixed-point theorem, we must verify that $\mathfrak{B}$ is upper semicontinuous. It suffices to check upper semicontinuity of each component $\mathfrak{B}_{i}$ separately [fan_fixed-point_1952, Lemma 3, p. 124]. Furthermore, by [yannelis_existence_1983, Lemma 6.1, p. 241], it is enough to show that each constraint map
${\bf X}_{i}:X_{-i}^{\rm ad}\to 2^{X_{i}^{\rm ad}}$
is upper semicontinuous. This follows from the fact that ${\bf X}_{i}$ has closed graph and that $X_{i}^{\rm ad}$ is compact: one verifies that $\left({\bf X}_{i}\right)^{-1}(C)$ is closed for every closed $C\subset X_{i}^{\rm ad}$.
All conditions required by Kakutani’s fixed-point theorem are therefore satisfied, so $\mathfrak{B}$ admits a fixed point in $X^{\rm ad}$. As argued above, such a point is a geometric generalized Nash equilibrium. ∎
The theorem above is the geometric counterpart of Theorem 3.4 for analytic games and highlights the role of lower semicontinuity of the constraint maps in the geometric setting. Just as in the analytic case, we introduce a Nikaido–Isoda type map $\Upsilon:X\times X\to 2^{X}$ by
$\Upsilon(x,y)=\prod_{i=1}^{N}\mathcal{P}_{i}(y_{i},x_{-i}).$ (7.2)
The next result is the geometric counterpart of Lemma 4.1.
Lemma 7.4.
Let $G=({\bf X},\mathcal{P})$ be a geometric game. A strategy bundle $\overline{x}\in X^{\rm ad}$ is a (geometric) Nash equilibrium of $G$ if and only if
$\overline{x}\in{\bf X}(\overline{x})\cap\overline{\Upsilon(\overline{x},y)}\quad\text{for all }y\in{\bf X}(\overline{x}).$
Proof.
Necessity follows directly from the definition of an equilibrium and from the construction of $\Upsilon$. Indeed, if $\overline{x}$ is a Nash equilibrium, then no player strictly prefers any other feasible strategy, so $\overline{x}$ must lie in the closure of all sets $\Upsilon(\overline{x},y)$ for $y\in{\bf X}(\overline{x})$.
For sufficiency, assume that $\overline{x}\in{\bf X}(\overline{x})$ and that
$\overline{x}\in\overline{\Upsilon(\overline{x},y)}\quad\text{for all }y\in{\bf X}(\overline{x}).$
Suppose, towards a contradiction, that $\overline{x}$ is not a Nash equilibrium. Then there exist $i\in I$ and $x_{i}\in{\bf X}_{i}(\overline{x}_{-i})$ such that $x_{i}\in\mathcal{P}_{i}(\overline{x})$; that is, player $i$ strictly prefers $x_{i}$ to $\overline{x}_{i}$. This implies $\overline{x}_{i}\notin\overline{\mathcal{P}_{i}(x_{i},\overline{x}_{-i})}.$ But since $(x_{i},\overline{x}_{-i})\in{\bf X}(\overline{x})$, the assumption $\overline{x}\in\overline{\Upsilon(\overline{x},(x_{i},\overline{x}_{-i}))}$ forces $\overline{x}_{i}\in\overline{\mathcal{P}_{i}(x_{i},\overline{x}_{-i})},$ and this is a contradiction. Therefore, $\overline{x}$ must be a Nash equilibrium. ∎
The notion of a variational equilibrium can also be formulated in geometric terms. We keep the name variational to stress the parallel to the analytic case, even though no variational inequality is written explicitly.
Definition 7.5 (Variational equilibrium).
Suppose that ${\bf X}_{i}$ is given by (2.2). A strategy bundle $\overline{x}\in\mathcal{C}$ is called a variational Nash equilibrium of a geometric game $G=({\bf X},\mathcal{P})$ if
$\overline{x}\in\mathcal{C}\cap\overline{\Upsilon(\overline{x},y)}\quad\text{for all }y\in\mathcal{C}.$
The following characterization in terms of the preference map $\mathcal{P}$ is often more convenient for existence arguments; see Theorem 7.7.
Lemma 7.6.
Let $G=({\bf X},\mathcal{P})$ be a geometric game and $\mathcal{C}\subset X$ closed. A strategy bundle $\overline{x}\in\mathcal{C}$ is a variational equilibrium if and only if
$\mathcal{P}(\overline{x})\cap\mathcal{C}=\emptyset.$
Proof.
By definition, $\overline{x}\in\mathcal{C}$ is a variational equilibrium if no strategy in $\mathcal{C}$ is strictly preferred to $\overline{x}$ by all players, that is, if there is no $y\in\mathcal{C}$ with $y\in\mathcal{P}(\overline{x})$. This is equivalent to the condition $\mathcal{P}(\overline{x})\cap\mathcal{C}=\emptyset.$
Theorem 7.7.
Let $\mathcal{C}$ be a compact and convex subset of $X$ such that $X^{\rm{ad}}\cap\mathcal{C}\neq\emptyset$ and assume that the structure (2.2) holds for the constraint maps. If $X$ is a separable Banach space, then every generalized geometric game has a variational equilibrium.
Proof.
The proof follows the same general idea as Theorem 7.3, so we only recall the main steps. Define
$F=\{x\in X^{\rm{ad}}\cap\mathcal{C}:\mathcal{P}(x)\cap\mathcal{C}\neq\emptyset\},\qquad F(x):=\mathcal{P}(x)\cap\mathcal{C}.$
On $F$, the map $F(\cdot)$ admits a continuous selection $f$. We then define $\mathfrak{B}(x)=f(x)$ for $x\in F$ and $\mathfrak{B}(x)=\mathcal{C}$ for $x\in X^{\rm{ad}}\setminus F$, and show that $\mathfrak{B}$ has a fixed point in $X^{\rm{ad}}$ which cannot belong to $F$. This fixed point is a variational equilibrium. ∎
Potential structures can also be introduced at the geometric level. The next definition parallels Definition 4.8.
Definition 7.8 (Potential game).
Assume that the constraint maps are of the form (2.2), that $\mathcal{C}\subset X$ is closed, and that $X$ is a separable Banach space. A geometric game $G=({\bf X},\mathcal{P})$ is called a potential game if there exists a continuous function $\mathcal{G}:X\to\mathbb{R}$ and a function $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ such that:
  • (i)
    $g(t_{k})\to 0$ implies $t_{k}\to 0$;
  • (ii)
    for all $i\in I$, all $x_{-i}\in X_{-i}$, and all $y_{i},z_{i}\in{\bf X}_{i}(x_{-i})$, the relation
    $z_{i}\in\mathcal{P}_{i}(y_{i},x_{-i})$
    implies
    $\mathcal{G}(y_{i},x_{-i})-\mathcal{G}(z_{i},x_{-i})\geqslant g\big(\|y_{i}-z_{i}\|_{X_{i}}\big).$
The function $\mathcal{G}$ is called a potential for the geometric game $G$.
The results in this section show that many analytic notions (Nikaido–Isoda maps, variational equilibria, potential games) have natural geometric counterparts. One could ask whether the graph-convex and KKM-based existence results from Section 5 also extend to the geometric setting. The authors believe that this should be possible, but such an extension would require:
  • a careful analysis of how graph-convexity or the KKM property for $\bf X$} interacts with the open-graph assumptions on the preference maps $\mathcal{P}_{i}$;
  • refined fixed-point theorems for set-valued maps on infinite-dimensional spaces that combine geometric properties of the feasible sets and of the preference correspondences;
  • an adaptation of the selection arguments used in Theorems 7.3 and 7.7 to situations where lower semicontinuity is not available.
Developing this additional machinery would significantly increase the technical length of the paper. For this reason, we chose to restrict the geometric part to a parallel of the lower semicontinuity-based theory and to highlight, rather than fully resolve, the questions that arise when one attempts to replace lower semicontinuity by purely geometric assumptions.
@@ 8 Conclusions and Outlook
On the geometric side, we showed that many analytic constructions have direct counterparts in terms of preference maps and that variational equilibria and potential structures can also be formulated without explicit objective functionals.

### v2

@@ 4 Multiplier bias and uniqueness of an equilibrium for jointly convex games
where ${\bf E}(\cdot)$ denotes the set of equilibria. This means that the set of Nash equilibria is unaffected by positive rescaling of the objective functions. An interesting aspect, however, is that except in the orthogonal case, i.e., when $\mathcal{C}$ has the form $\mathcal{C}=\mathcal{C}_{1}\times\cdots\times\mathcal{C}_{N}$, we have ${\bf VE}(G_{r})\neq{\bf VE}(G)$, where ${\bf VE}(\cdot)$ is the set of variational equilibria. We show below, extending the results of [32] to infinite dimensions, that an $r$-dependent structural condition on $G_{r}$ ensures that ${\bf VE}(G_{r})$ is a singleton. Hence, any numerical method designed to compute a variational equilibrium will converge to the same one. Since ${\bf VE}(G_{r})\subset{\bf E}(G_{r})={\bf E}(G)$ and $r\in(\mathbb{R}_{+}^{*})^{N}$ must be chosen so that the structural condition holds, we are still computing a Nash equilibrium of the original game.
@@
We begin with the non-generalized situation, i.e., we assume that ${\bf X}(x)\equiv X^{\rm{ad}}$ even if ${\bf X}_{i}$ has the form (2.2). This corresponds to the case in which $\mathcal{C}$ is a Cartesian product. The result below extends [32, Theorem 2] to the infinite-dimensional setting. In its proof we need the normal cone to $\mathcal{C}$ at some $x\in\mathcal{C}$ which is given by
@@
From convex optimization along with the Nikaido-Isoda function formulation of a Nash equilibrium, it follows that if $\overline{x}\in\mathcal{C}$ is a Nash equilibrium, then there exists $\mu_{\overline{x}}\in N_{\mathcal{C}}(\overline{x})$ such that
@@
$\alpha=d(\overline{x},r)(\overline{y}-\overline{x})+d(\overline{y},r)(\overline{x}-\overline{y}),$
$\beta=\mu_{\overline{x}}(r\cdot(\overline{y}-\overline{x}))+\mu_{\overline{y}}(r\cdot(\overline{x}-\overline{y})).$
@@
It follows from 4.5 and 4.6 that when the constraint map is constant, the variational equilibrium is unique and independent of $r$. However, in the non-constant case, although the scaling vector $r$ cannot alter the set of Nash equilibria, it may affect the set of variational equilibria – a feature of practical interest given their computational advantages over the full equilibrium set. The next proposition quantifies this effect.
@@
We do not include the proofs of 4.6 and 4.7 because the arguments follow the same structure as 4.5. The corresponding finite-dimensional proofs are found in [32].
@@
Diagonal strict convexity also implies a strong stability property, naturally formulated in terms of an associated ordinary differential equation type dynamical system that generates a trajectory of strategies. This dynamical perspective has served as a foundation for algorithms computing variational equilibria and offers a complementary viewpoint to modern numerical methods for variational inequalities; see, for example, [11, 12] and references therein. Beyond its algorithmic relevance, this interpretation provides a useful conceptual and historical connection between equilibrium theory and dynamical systems.